OpenStudy (anonymous):

match with the correct solution log b 121 =2

OpenStudy (anonymous):

here are the answer choices 3 27 81 9 11 243

OpenStudy (anonymous):

the b is below log

OpenStudy (anonymous):

its at the top

OpenStudy (anonymous):

write this out in exponential form...

OpenStudy (anonymous):

\[\huge b^2 = 121\] What can be b?? Take the square root..

OpenStudy (anonymous):

\[\huge log_b121=2\Leftrightarrow b^2=121 \]

OpenStudy (anonymous):

Which number has the square 121 ???

OpenStudy (anonymous):

yall are confusing me

OpenStudy (anonymous):

What you are not getting ??

OpenStudy (anonymous):

On previous one I have given you a formula you remember that???

OpenStudy (anonymous):

idk but its ohk i kinda understand i think is it 9?

OpenStudy (anonymous):

is 9^2 = 121 ???

OpenStudy (anonymous):

idk just forget it no its 81

OpenStudy (anonymous):

\[\huge \color{green}{\log_y(x) = b \implies y^b = x}\] Here, y = b b = 2 and x = 121

OpenStudy (anonymous):

11!

OpenStudy (anonymous):

Yes you are right now..

OpenStudy (anonymous):

11^2=121!

OpenStudy (anonymous):

lol oh wow im dumb

OpenStudy (anonymous):

Just practice and work hard.. You will surely get this..

OpenStudy (anonymous):

thanks!